OpenStudy (anonymous):

identify the following curves.give in standard form: 1)16y^2-400+25x^2 2)16x^2-144=9y^2

OpenStudy (anonymous):

solve for y

OpenStudy (mertsj):

\[25x^2+16y^2=400\] \[\frac{x^2}{16}+\frac{y^2}{25}=1\]

OpenStudy (mertsj):

Ellipse

OpenStudy (mertsj):

\[16x^2-9y^2=144\] \[\frac{x^2}{9}-\frac{y^2}{16}=1\]

OpenStudy (mertsj):

Hyperbola

OpenStudy (anonymous):

is it the standard form y=mx+c ?